OpenStudy (anonymous):

Keep the pattern going 2% .07 3/25 17% 22% _____

OpenStudy (anonymous):

27%, or .27, or 27/100

OpenStudy (diyadiya):

2%= 2/100=0.02 3/25=0.12 17%=17/100=0.17 22%=22/100=0.22 So this is 0.02, 0.07 , 0.12 ,0.17 ,0.22 ** 0.07-0.02=0.05 0.12-0.07=0.05 so the common difference is 0.05 so add 0.05 to 0.22 0.22+0.05 =0.27 = 27%